OpenStudy (anonymous):

find the unit rate for each situation round to the nearest hundredth if neccesarry reading 53 pages in 2 hr and 592 students in 17 classrooms

OpenStudy (anonymous):

53 / 2 = 26.5 pages per hour. 592 / 17 = 34.82 students each classroom.

jimthompson5910 (jim_thompson5910):

53/2 = 26.5 pages per hour -------------------------- 592/17 = 34.82 students per classroom